# The completeness problem on 3-dimensional non-unimodular Lie groups

### v2

@@ 4.1.2. Incomplete geodesic fields with no idempotents
Their energy is as a first integral and, moreover, $y$ (and so $y^{2}$) is clearly another first integral. Therefore, the linear combinations
$\begin{array}[]{lclcl}\mathscr{I}_{4,0}&=&\tfrac{1}{2}(x^{2}+y^{2}+2yz)+\tfrac{1}{2}y^{2}&=&\tfrac{1}{2}x^{2}+y^{2}+yz\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \mathscr{I}_{5,0}&=&\tfrac{1}{2}(x^{2}-y^{2}+2yz)-\tfrac{1}{2}y^{2}&=&\tfrac{1}{2}x^{2}-y^{2}+yz\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \mathscr{I}_{9}&=&\tfrac{1}{2}(x^{2}+2yz)&=&\tfrac{1}{2}x^{2}+yz\\ \end{array}$
@@ 4.2.1. Complete geodesic fields whose integral curves are all bounded
The Lorentzian metrics for which $\mathfrak{z}$ is timelike are $\mathscr{Q}_{1}$, $\mathscr{Q}_{4,r}$ $(r<0)$, $\mathscr{Q}_{5,s}$ $(-1<s<0)$ and $\mathscr{Q}_{8}$. Their geodesics fields are given by
$\begin{array}[]{ll}\mathscr{F}_{1}:\begin{cases}\dot{x}=-y^{2}\\ \dot{y}=xy\\ \dot{z}=0\end{cases}&\mathscr{F}_{4,r}:\begin{cases}\dot{x}=-y^{2}-yz\\ \dot{y}=\frac{rx}{r-1}(y+z)\\ \dot{z}=\frac{-x}{r-1}(y+z)\end{cases}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \mathscr{F}_{5,s}:\begin{cases}\dot{x}=y^{2}-yz\\ \dot{y}=\frac{sx}{s+1}(y-z)\\ \dot{z}=\frac{x}{s+1}(z-y)\end{cases}&\mathscr{F}_{8}:\begin{cases}\dot{x}=-yz\\ \dot{y}=xz\\ \dot{z}=xz\end{cases}\end{array}$
@@
[…]
and observe that they are all positive definite. Thus, the geodesic fields $\mathscr{F}_{1}$, $\mathscr{F}_{4,r}$ $(r<0)$, $\mathscr{F}_{5,s}$ $(-1<s<0)$, $\mathscr{F}_{8}$ have all their integral curves are bounded and are, therefore, complete.
@@
It is well-known by [RS94, Prop. 2.1] that the existence of a timelike Killing field on a Lorentzian manifold with some boundedness conditions implies geodesic completeness. In our case, the Lie algebra vector $e_{3}$ is central and this implies that the corresponding left-invariant vector field $E_{3}$ is also right-invariant. In particular, $E_{3}$ is a Killing field which has constant energy. When $E_{3}$ is timelike, completeness of our geodesic fields then follows immediately. Moreover, it is worth noting that our linear first integrals were expected. The fact that $E_{3}$ is Killing implies that every geodesic field has a linear first integral given by $q(e_{3},-)$, where $q$ is the metric in question.
4.2.2. Complete geodesic fields whose integral curves are all unbounded
The remaining Lorentzian metric in our discussion is the only one for which $\mathfrak{z}$ is lightlike and orthogonal to $\mathfrak{d}$, that is, $\mathscr{Q}_{12}$. Its geodesic field is given by
@@ 4.2.2. Complete geodesic fields whose integral curves are all unbounded
Thus, $\mathscr{F}_{12}$ is a complete geodesic field and its integral curves are all unbounded.
@@
5. (In)completeness on $\mathfrak{h}(\lambda)$, $10<|\lambda|<$ 1: proof of Theorem 1.2
In this section, we classify Lorentzian metrics on the Lie algebras $\mathfrak{h}(\lambda)$, $10<|\lambda|<$ 1, with respect to their geodesic (in)completeness and give the proof of Theorem 1.2. Our strategy here will be similar to that of the Sec. 4.
Inspecting Table 1 once more, we see that the family $\mathscr{Q}_{5,s}$, with $-1<s<0$, corresponds to the metric normal forms for which both $e_{2}$ and $e_{3}$ are timelike. On the other hand, $\mathscr{Q}_{8}$ is the only one for which $e_{3}$ is timelike and $e_{2}$ is lightlike. We start by showing that every other metric normal form is incomplete and proceed with the analysis of the completeness of $\mathscr{Q}_{5,s}$, $-1<s<0$, and $\mathscr{Q}_{8}$.
@@ 5.1.1. Incomplete geodesic fields with idempotents
  • $\mathscr{F}_{2}$: $(1,1,0),(1,-1,0)$;
  • $\mathscr{F}_{3}$: $(1,1,0),(1,-1,0)$, $\left(\frac{1}{\lambda},0,\frac{1}{\lambda}\right)$, $\left(\frac{1}{\lambda},0,\frac{-1}{\lambda}\right)$;
@@ 5.1.2. Incomplete geodesic fields with no idempotents
which, in turn, implies the existence of incomplete integral curves of $\mathscr{F}_{5,0}$ by [CFZ24, Lemma 2.3], since $\frac{1-\lambda}{2}y^{2}$ is nonnegative for $10<|\lambda|<$ 1. Thus, $\mathscr{F}_{5,0}$ is incomplete.
@@ 5.2.1. Complete geodesic fields whose integral curves are not all bounded
Since $\{z=0\}$ is invariant, then so are the half spaces $H^{+}=\{z>0\}$ and $H^{-}=\{z<0\}$. It can be easily checked that, besides the energy, no other quadratic first integrals exist. However, we have the following.
@@ 5.2.2. Complete geodesic fields whose integral curves are all bounded
The remaining Lorentzian metrics in our analysis are the ones in the family $\mathscr{Q}_{5,s}$, $-1<s<0$. As seen from Table 1, the geodesic field $\mathscr{F}_{5,s}$ looks somewhat complicated. However the existence of the two invariant planes $\{y=z\}$ and $\{y=-sz\}$ hints at the following linear change of variables
@@
and, indeed, it becomes simpler. The energy is given by $e_{s}(x,\xi,\eta)=x^{2}+\frac{1}{s+1}(\xi^{2}+2\xi\eta-s\eta^{2})$. The invariant plane $\{\eta=0\}$ is spacelike and all integral curves with initial condition $(x_{o},\xi_{o},\eta_{o})$ with $\eta_{o}=0$ and bounded and, hence, complete. The half-spaces $H^{+}=\{\eta>0\}$ and $H^{-}=\{\eta<0\}$ are also invariant. Moreover, we have the following partially defined first integral.
@@
$-\frac{s}{s+1}\eta^{2}\left(1-\frac{c^{2}}{s}\eta^{2(\lambda-1)}-\frac{2c}{s}\eta^{\lambda-1}\right)\leq k.$
It is important to observe here that not only $\lambda-1<0$ but also that $-\frac{s}{s+1}>0$ since $-1<s<0$.
Thus, $\eta$ cannot go to infinity and is, therefore, bounded. The rest of the argument to prove that $\mathscr{F}_{5,s}$ is a complete vector fields is now analogous to that of Subsec. 5.2.1.
@@ 6. The limiting Lie algebras $\mathfrak{h}(\lambda)$, with $\lambda=\pm 1$
The geodesic fields of Table 1 in Sec. 3 clearly extend to the limiting cases $\lambda=\pm 1$ (although care should be taken concerning the existence of idempotents). Together with the metric normal forms of Rem. 2.1, we can quickly solve the completeness problem for $\mathfrak{h}(1)$ and $\mathfrak{h}(-1)$, recovering the result of [Gue96] in dimension 3 and correcting the statement in [BM08, Prop. 3].
@@ 6.1. Case $\lambda=1$
The Lorentzian normal forms are $\mathscr{Q}_{1},\mathscr{Q}_{3}$ and $\mathscr{Q}_{11}$ and they are all incomplete with idempotents.
@@ 6.2. Case $\lambda=-1$
From Table 1, we readily check that all metric normal forms given in Rem. 2.1 are incomplete with idempotents except for $\mathscr{Q}_{5,s}$ ($-1<s\leq 0$) and $\mathscr{Q}_{9}$. We now show that these remaining metrics are complete. The geodesic field associated to $\mathscr{Q}_{9}$ is given by
@@
As for the family $\mathscr{Q}_{5,s}$ ($-1<s\leq 0$), we can easily check that, besides the energy $e(x,y,z)=x^{2}-y^{2}+2yz+sz^{2}$, the first integral of Subsec. 5.2.2 extends to $\lambda=-1$ and yields, in fact, the quadratic first integral $f(x,y,z)=y^{2}+(s-1)yz-sz^{2}$. Taking $e(x,y,z)+2f(x,y,z)=x^{2}+(y+sz)^{2}-(s^{2}+s)z^{2}$, we obtain a positive definite quadratic first integral, for all $-1<s<0$. So, all integral curves of $\mathscr{F}_{5,s}$ ($-1<s<0$) are bounded. For $s=0$, we have the geodesic field
@@
In summary, indeed a Lorentzian metric is incomplete if and only it has an idempotent. However, not all complete metrics admit a positive definite quadratic first integral, as is the case of $\mathscr{Q}_{9}$ and $\mathscr{Q}_{5,0}$.
@@ 7. Incompleteness of 3-dimensional non-unimodular Lie algebras
In this section, we wish to finalize the proof that every non-unimodular 3-dimensional Lie algebra admits an incomplete Lorentzian metric. For the Lie algebra $\mathfrak{h}(1)$, this was done in [Gue96], for $\mathfrak{psh}$ in [CFZ24] and for $\mathfrak{h}(\lambda)$, $|\lambda|<1$, in the present paper. It thus remains to show that, for every $\mu>0$, the Lie algebra $\mathfrak{e}(\mu)$ admits at least one incomplete metric.
@@ Acknowledgments
The authors acknowledge the support of CMAT (Centro de Matemática da Universidade do Minho). Their research was financed by Portuguese Funds through FCT (Fundação para a Ciência e a Tecnologia, I.P.) within the Projects UIDB/00013/2020, UIDP/00013/2020, UID/00013: Centro de Matemática da Universidade do Minho (CMAT/UM) and also within the doctoral grant UI/BD/154255/2022 of the first named author.

### v3

@@ 5.1.2. Incomplete geodesic fields with no idempotents
Their energy is a first integral and, moreover, $y$ (and so $y^{2}$) is clearly another first integral. Therefore, the linear combinations
[…]
@@ 5.2.1. Complete geodesic fields whose integral curves are all bounded
The Lorentzian metrics for which $\mathfrak{z}$ is timelike are $\mathscr{Q}_{1}$, $\mathscr{Q}_{4,r}$ $(r<0)$, $\mathscr{Q}_{5,s}$ (1<s<0)(-1<s<0) and $\mathscr{Q}_{8}$. Their geodesic fields are given by
$\begin{array}[]{ll}\mathscr{F}_{1}:\begin{cases}\dot{x}=-y^{2}\\ \dot{y}=xy\\ \dot{z}=0\end{cases}\qquad&\mathscr{F}_{4,r}:\begin{cases}\dot{x}=-y^{2}-yz\\ \dot{y}=\frac{rx}{r-1}(y+z)\\ \dot{z}=\frac{-x}{r-1}(y+z)\end{cases}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \mathscr{F}_{5,s}:\begin{cases}\dot{x}=y^{2}-yz\\ \dot{y}=\frac{sx}{s+1}(y-z)\\ \dot{z}=\frac{x}{s+1}(z-y)\end{cases}&\mathscr{F}_{8}:\begin{cases}\dot{x}=-yz\\ \dot{y}=xz\\ \dot{z}=xz\end{cases}\end{array}$
@@
$\begin{array}[]{lclcl}\mathscr{I}_{1}&=&(x^{2}+y^{2}-z^{2})+2z^{2}&=&x^{2}+y^{2}+z^{2}\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \mathscr{I}_{4,r}&=&(x^{2}+y^{2}+rz^{2}+2yz)-\frac{r-2}{r(r-1)}(y+rz)^{2}&=&x^{2}+\frac{1}{r(r-1)}((r-1)^{2}y^{2}+(y+rz)^{2})\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \mathscr{I}_{5,s}&=&(x^{2}-y^{2}+sz^{2}+2yz)-\frac{s+2}{s(s+1)}(y+sz)^{2}&=&x^{2}-\frac{1}{s(s+1)}((s+1)^{2}y^{2}+(y+sz)^{2})\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \mathscr{I}_{8}&=&(x^{2}+2yz-z^{2})+2(y-z)^{2}&=&x^{2}+y^{2}+(y-z)^{2}\end{array}$
and observe that they are all positive-definite. Thus, the geodesic fields $\mathscr{F}_{1}$, $\mathscr{F}_{4,r}$ $(r<0)$, $\mathscr{F}_{5,s}$ (1<s<0)(-1<s<0), $\mathscr{F}_{8}$ have all their integral curves bounded and are, therefore, complete.
@@
It is well known that the existence of a timelike Killing field on a Lorentzian manifold with some boundedness conditions implies geodesic completeness, [RS94, Prop. 2.1]. In our case, the Lie algebra vector $e_{3}$ is central and this implies that the corresponding left-invariant vector field $E_{3}$ is also right-invariant. In particular, $E_{3}$ is a Killing field which has constant energy. When $E_{3}$ is timelike, completeness of our geodesic fields then follows immediately. Moreover, it is worth noting that our linear first integrals were expected. The fact that $E_{3}$ is Killing implies that every geodesic field has a linear first integral given by $q(e_{3},-)$, where $q$ is the metric in question.
5.2.2. Complete geodesic fields whose non-stationary integral curves are all unbounded
The remaining Lorentzian metric in our discussion is the only case where $\mathfrak{z}$ is lightlike and orthogonal to $\mathfrak{d}$, that is, $\mathscr{Q}_{12}$. Its geodesic field is given by
@@ 5.2.2. Complete geodesic fields whose non-stationary integral curves are all unbounded
Thus, $\mathscr{F}_{12}$ is a complete geodesic field and its non-stationary integral curves are all unbounded.
@@
6. (In)completeness on $\mathfrak{h}(\lambda)$, $0<|\lambda|<1$: proof of Theorem 1.2
In this section, we classify Lorentzian metrics on the Lie algebras $\mathfrak{h}(\lambda)$, $0<|\lambda|<1$, with respect to their geodesic (in)completeness and give the proof of Theorem 1.2. Our strategy here will be similar to that of Sec. 5.
Inspecting Table 1 once more, we see that the family $\mathscr{Q}_{5,s}$, with 1<s<0-1<s<0, corresponds to the metric normal forms for which both $e_{2}$ and $e_{3}$ are timelike. On the other hand, $\mathscr{Q}_{8}$ is the only one for which $e_{3}$ is timelike and $e_{2}$ is lightlike. We start by showing that every other metric normal form is incomplete and proceed with the analysis of the completeness of $\mathscr{Q}_{5,s}$, 1<s<0-1<s<0, and $\mathscr{Q}_{8}$.
@@ 6.1.1. Incomplete geodesic fields with idempotents
  • $\mathscr{F}_{2}$: (1,1,0),(1,1,0)(1,1,0),(1,-1,0);
  • $\mathscr{F}_{3}$: (1,1,0),(1,1,0)(1,1,0),(1,-1,0), $\left(\frac{1}{\lambda},0,\frac{1}{\lambda}\right)$, $\left(\frac{1}{\lambda},0,\frac{-1}{\lambda}\right)$;
@@ 6.1.2. Incomplete geodesic fields with no idempotents
which, in turn, implies the existence of incomplete integral curves of $\mathscr{F}_{5,0}$ by Lem. 2.1, since $\frac{1-\lambda}{2}y^{2}$ is nonnegative for $0<|\lambda|<1$. Thus, $\mathscr{F}_{5,0}$ is incomplete.
@@ 6.2.1. Complete geodesic fields whose integral curves are not all bounded
Since $\{z=0\}$ is invariant, then so are the half-spaces $H^{+}=\{z>0\}$ and $H^{-}=\{z<0\}$. It can be easily checked that, besides the energy, no other quadratic first integrals exist. However, we have the following.
@@ 6.2.2. Complete geodesic fields whose integral curves are all bounded
The remaining Lorentzian metrics in our analysis are the ones in the family $\mathscr{Q}_{5,s}$, 1<s<0-1<s<0. As seen from Table 1, the geodesic field $\mathscr{F}_{5,s}$ looks somewhat complicated. However, the existence of the two invariant planes $\{y=z\}$ and $\{y=-sz\}$ hints at the following linear change of variables
@@
and, indeed, it becomes simpler. The energy is given by $e_{s}(x,\xi,\eta)=x^{2}+\frac{1}{s+1}(\xi^{2}+2\xi\eta-s\eta^{2})$. The invariant plane $\{\eta=0\}$ is spacelike and all integral curves with initial condition $(x_{o},\xi_{o},\eta_{o})$ with $\eta_{o}=0$ are bounded, and hence complete. The half-spaces $H^{+}=\{\eta>0\}$ and $H^{-}=\{\eta<0\}$ are also invariant. Moreover, we have the following partially defined first integral.
@@
$\frac{s}{s+1}\eta^{2}\left(1-\frac{c^{2}}{s}\eta^{2(\lambda-1)}-\frac{2c}{s}\eta^{\lambda-1}\right)\leq k.$
It is important to observe here that not only $\lambda-1<0$ but also that $\frac{s}{s+1}>0$ since 1<s<0-1<s<0.
Thus, $\eta$ cannot go to infinity and is, therefore, bounded. The rest of the argument to prove that $\mathscr{F}_{5,s}$ is a complete vector field is now analogous to that of Subsec. 6.2.1.
@@ 7. The limiting Lie algebras $\mathfrak{h}(\lambda)$, with $\lambda=\pm 1$
The geodesic fields of Table 1 in Sec. 4 clearly extend to the limiting cases $\lambda=\pm 1$ (although care should be taken concerning the existence of idempotents). Together with the metric normal forms of Rem. 3.1, we can quickly solve the completeness problem for $\mathfrak{h}(1)$ and $\mathfrak{h}(-1)$, recovering the result of [Gue96] in dimension 3 and correcting the statement in [BM08, Prop. 3].
@@ 7.1. Case $\lambda=1$
The Lorentzian normal forms are $\mathscr{Q}_{1},\mathscr{Q}_{3}$, and $\mathscr{Q}_{11}$, and they are all incomplete with idempotents.
@@ 7.2. Case $\lambda=-1$
From Table 1, we readily check that all metric normal forms given in Rem. 3.1 are incomplete with idempotents except for $\mathscr{Q}_{5,s}$ ($1<s\leq 0$) and $\mathscr{Q}_{9}$. We now show that these remaining metrics are complete. The geodesic field associated to $\mathscr{Q}_{9}$ is given by
@@
As for the family $\mathscr{Q}_{5,s}$ ($1<s\leq 0$), we can easily check that, besides the energy $e(x,y,z)=x^{2}-y^{2}+2yz+sz^{2}$, the first integral of Subsec. 6.2.2 extends to $\lambda=-1$ and yields, in fact, the quadratic first integral $f(x,y,z)=y^{2}+(s-1)yz-sz^{2}$. Taking $e(x,y,z)+2f(x,y,z)=x^{2}+(y+sz)^{2}-(s^{2}+s)z^{2}$, we obtain a positive-definite quadratic first integral, for all 1<s<0-1<s<0. So, all integral curves of $\mathscr{F}_{5,s}$ (1<s<0-1<s<0) are bounded. For $s=0$, we have the geodesic field
@@
In summary, indeed a Lorentzian metric is incomplete if and only if it has an idempotent. However, not all complete metrics admit a positive-definite quadratic first integral, as is the case of $\mathscr{Q}_{9}$ and $\mathscr{Q}_{5,0}$.
@@ 8. Incompleteness of 3-dimensional non-unimodular Lie algebras
In this section, we wish to finalize the proof that every non-unimodular 3-dimensional Lie algebra admits an incomplete Lorentzian metric. For the Lie algebra $\mathfrak{h}(1)$, this was done in [Gue96]; for $\mathfrak{psh}$ in [CFZ25a]; and for $\mathfrak{h}(\lambda)$, $|\lambda|<1$, in the present paper. It thus remains to show that, for every $\mu>0$, the Lie algebra $\mathfrak{e}(\mu)$ admits at least one incomplete metric.
@@ Acknowledgments
The authors acknowledge the support of CMAT (Centro de Matemática da Universidade do Minho). Their research was financed by Portuguese Funds through FCT (Fundação para a Ciência e a Tecnologia, I.P.) within the projects UIDB/00013/2020, UIDP/00013/2020, UID/00013/2025 and also within the doctoral grant UI/BD/154255/2022 of the first named author.